### v3

# EVODiff: Entropy-aware Variance Optimized Diffusion Inference

Abstract
Diffusion models (DMs) excel in image generation but suffer from slow inference and training-inference discrepancies. Although gradient-based solvers for DMs accelerate denoising inference, they often lack theoretical foundations in information transmission efficiency. In this work, we introduce an information-theoretic perspective on the inference processes of DMs, revealing that successful denoising fundamentally reduces conditional entropy in reverse transitions. This principle leads to our key insights into the inference processes: (1) data prediction parameterization outperforms its noise counterpart, and (2) optimizing conditional variance offers a reference-free way to minimize both transition and reconstruction errors. Based on these insights, we propose an entropy-aware variance optimized method for the generative process of DMs, called EVODiff, which systematically reduces uncertainty by optimizing conditional entropy during denoising. Extensive experiments on DMs validate our insights and demonstrate that our method significantly and consistently outperforms state-of-the-art (SOTA) gradient-based solvers. For example, compared to the DPM-Solver++, EVODiff reduces the reconstruction error by up to 45.5% (FID improves from 5.10 to 2.78) at 10 function evaluations (NFE) on CIFAR-10, cuts the NFE cost by 25% (from 20 to 15 NFE) for high-quality samples on ImageNet-256, and improves text-to-image generation while reducing artifacts. Code is available at https://github.com/ShiguiLi/EVODiff.
1 Introduction
Diffusion models (DMs) [1, 2, 3] have emerged as powerful generative models, achieving success in tasks such as image synthesis and editing [4, 5], text-to-image generation [6], voice synthesis [7], and video generation [8]. DMs generate high-fidelity samples by simulating a denoising process that iteratively refines noisy inputs through diffusion inference guided by a trained model. The model is trained via a forward process that corrupts data with Gaussian noise across multiple scales.
Despite their impressive performance, DMs face the challenge of a slow refinement process and a discrepancy between training and inference [2, 9, 10]. To address this, training-free inference methods reformulate the denoising process as the solution to an ODE using numerical techniques. Such examples include PNDM [11], EDM [12], DPM-Solver [13], DEIS [14], SciRE-Solver [15], UniPC [16], and DPM-Solver-v3 [17]. Despite their empirical success, these ODE-focused methods lack an information-theoretic foundation. A central limitation is their neglect of information transmission efficiency. This theoretical gap suggests that the principles governing inference remain underexplored.
DDIM
EVODiff
Figure 1: Illustration of conditional entropy reduction during diffusion model inference. Our EVODiff (blue) achieves lower conditional entropy in reverse transitions compared to DDIM (gray).
Our work addresses this gap by developing an information-theoretic framework for diffusion inference centered on conditional entropy dynamics. In this view, the forward diffusion process systematically increases conditional entropy as noise is added, while the reverse process seeks to recover lost information through denoising. Unlike gradient-based ODE solvers that primarily focus on numerical approximation, our framework reveals that effective denoising fundamentally operates by reducing conditional entropy during reverse transitions, a principle largely overlooked by existing methods. This insight not only guides algorithm design, but also offers a unified theoretical explanation of the varying inference efficiency across successful strategies, rooted in their entropy reduction efficiency.
Building on this framework, we propose EVODiff, an entropy-aware method that reduces conditional entropy by optimizing the conditional variance of each denoising iteration. Our approach provides three key technical advantages: (1) it enhances information transmission between successive denoising steps through entropy-reduction optimization; (2) it accelerates convergence by steering samples towards high-probability regions of the data distribution, drawing on principles from statistical physics [18, 19]; (3) it minimizes both transition errors and reconstruction errors through reference-free variance optimization (detailed in Section 3). These advantages lead to our main contributions:
  • We introduce an information-theoretic framework for diffusion inference, demonstrating that gradient-based methods enhance inference by reducing conditional entropy. Our analysis provides the first theoretical evidence for why data prediction parameterization outperforms its noise counterpart, theoretically grounding previous empirical findings [13, 20].
  • Guided by our insights, we propose EVODiff, an entropy-aware variance optimized diffusion inference method. Fundamentally, it differs from existing ODE-based approaches by directly targeting information recovery, not just approximating an ODE trajectory. This approach reduces both transition and reconstruction errors via principled variance optimization.
  • Extensive experiments validate our insights and demonstrate significant improvements in inference. EVODiff reduces FID by 45.5% on CIFAR-10 (from 5.10 to 2.78 at 10 NFE) and 45.5% on LSUN-Bedrooms (from 13.97 to 7.91 at 5 NFE) over strong solvers like DPM-Solver++ and UniPC, while also mitigating visual artifacts in text-to-image generation.
2 Background
Let $d$ denote the dimension of the data. The forward process of DMs defines a Markov sequence $\left\{\boldsymbol{x}_{t},t\in[0,T]\right\}$, where $\boldsymbol{x}_{0}\in\mathbb{R}^{d}$ is the starting state drawn from the data distribution $q(\boldsymbol{x}_{0})$ [1, 2]. This sequence is pushed forward with the transition kernel: $q\left(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{0}\right)=\mathcal{N}(\boldsymbol{x}_{t};\alpha_{t}\boldsymbol{x}_{0},\sigma_{t}^{2}\boldsymbol{I}),$ where $\alpha_{t}$ and $\sigma_{t}$ are the noise schedules and $\alpha_{t}^{2}/\sigma_{t}^{2}$ is the signal-to-noise ratio (SNR). This transition kernel is reformulated as the stochastic differential equation (SDE) [3]: $\mathrm{d}\boldsymbol{x}_{t}=f(t)\boldsymbol{x}_{t}\mathrm{~d}t+g(t)\mathrm{d}\boldsymbol{\omega}_{t},~\boldsymbol{x}_{0}\sim q\left(\boldsymbol{x}_{0}\right)$, where $\boldsymbol{\omega}_{t}$ denotes a Wiener process, $f(t):=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{~d}t},~g^{2}(t):=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{~d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{~d}t}\sigma_{t}^{2}$ [21]. In the denoising inference process, the reverse-time SDE of the forward diffusion process takes the form:
$\mathrm{d}\boldsymbol{x}_{t}=\left[f(t)\boldsymbol{x}_{t}-g^{2}(t)\nabla_{\boldsymbol{x}}\log q\left(\boldsymbol{x}_{t}\right)\right]\mathrm{d}t+g(t)\mathrm{d}\overline{\boldsymbol{\omega}}_{t},$ (1)
where $\overline{\boldsymbol{\omega}}_{t}$ represents a Wiener process. The inference generative process based on diffusion (or probability flow) ordinary differential equation (ODE) [3] is governed by $\mathrm{d}\boldsymbol{x}_{t}=\big(f(t)\boldsymbol{x}_{t}-\frac{1}{2}g^{2}(t)\nabla_{\boldsymbol{x}}\log q\left(\boldsymbol{x}_{t}\right)\big)\mathrm{~d}t$, where the marginal distribution $q\left(\boldsymbol{x}_{t}\right)$ of $\boldsymbol{x}_{t}$ is equivalent to that of $\boldsymbol{x}_{t}$ in the SDE of Eq. (1). The model is generally trained by minimizing the mean squared error (MSE) [2]:
$\mathbb{E}_{\boldsymbol{x}_{0},~\boldsymbol{\epsilon},~t}[w(t)\left\|\boldsymbol{\epsilon}_{\theta}\left(\alpha_{t}\boldsymbol{x}_{0}+\sigma_{t}\boldsymbol{\epsilon},t\right)-\boldsymbol{\epsilon}\right\|_{2}^{2}],$ (2)
where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$, $\boldsymbol{x}_{0}\sim q\left(\boldsymbol{x}_{0}\right)$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})$, $t\sim\mathcal{U}(0,T)$, and $w(t)$ is a weight function w.r.t. $t$.
Diffusion ODE. Based on the relationship of $\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t},t\right)=-\sigma_{t}\nabla_{\boldsymbol{x}}\log q\left(\boldsymbol{x}_{t}\right)$ [3], samples can be generated by the diffusion inference process from $T$ to 0 defined diffusion ODE:
$\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{~d}t}=f(t)\boldsymbol{x}_{t}+\frac{g^{2}(t)}{2\sigma_{t}}\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t},t\right),~\boldsymbol{x}_{T}\sim\mathcal{N}\left(\boldsymbol{0},\hat{\sigma}^{2}\boldsymbol{I}\right).$ (3)
By $\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t},t\right):=(\boldsymbol{x}_{t}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t},t\right))/\alpha_{t}$ [21], the data prediction ODE can be expressed as follows
$\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{~d}t}=\big(f(t)+\frac{g^{2}(t)}{2\sigma_{t}^{2}}\big)\boldsymbol{x}_{t}-\alpha_{t}\frac{g^{2}(t)}{2\sigma_{t}^{2}}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t},t\right).$ (4)
Remark 2.1.
A unified solution formula for both ODE formulations in (3) and (4) can be expressed as
$\boldsymbol{f}(\boldsymbol{x}_{t})-\boldsymbol{f}(\boldsymbol{x}_{s})=\int_{\kappa(s)}^{\kappa(t)}\boldsymbol{d}_{\theta}\left(\boldsymbol{x}_{\psi(\tau)},\psi(\tau)\right)\mathrm{d}\tau.$ (5)
where $\psi\left(\kappa(t)\right)=t$. When using noise prediction, we have $\boldsymbol{d}_{\theta}=\boldsymbol{\epsilon}_{\theta}$, $\boldsymbol{f}(\boldsymbol{x}_{t}):=\frac{\boldsymbol{x}_{t}}{\alpha_{t}}$ and $\kappa(t):=\frac{\sigma_{t}}{\alpha_{t}}$; when using data prediction, we have $\boldsymbol{d}_{\theta}=\boldsymbol{x}_{\theta}$, $\boldsymbol{f}(\boldsymbol{x}_{t}):=\frac{\boldsymbol{x}_{t}}{\sigma_{t}}$ and $\kappa(t):=\frac{\alpha_{t}}{\sigma_{t}}$ [22, 15].
Gradient-based Inference. Denote $h_{t_{i}}:=\kappa(t_{i-1})-\kappa(t_{i})$ and $\boldsymbol{\iota}(\boldsymbol{x}_{t_{i-1}}):=\boldsymbol{f}(\boldsymbol{x}_{t_{i-1}})-\boldsymbol{f}(\boldsymbol{x}_{t_{i}})$. By substituting the Taylor expansion of $\boldsymbol{d}_{\theta}\big(\boldsymbol{x}_{t_{i-1}},t_{i-1}\big)$ at $\tau_{t_{i}}$ into Eq. (5), we can obtain $\boldsymbol{\iota}(\boldsymbol{x}_{t_{i-1}})=\sum_{k=0}^{n}\frac{h_{t_{i}}^{k+1}}{(k+1)!}\boldsymbol{d}_{\theta}^{(k)}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)+\mathcal{O}(h_{t_{i}}^{n+2})$, where $\boldsymbol{d}_{\theta}^{(k)}\left(\boldsymbol{x}_{\psi(\tau)},\psi(\tau)\right):=\frac{\mathrm{d}^{k}\boldsymbol{d}_{\theta}\left(\boldsymbol{x}_{\psi(\tau)},\psi(\tau)\right)}{\mathrm{d}~\tau^{k}}$ as $k$-th order total derivative of $\boldsymbol{d}_{\theta}\left(\boldsymbol{x}_{\psi(\tau)},\psi(\tau)\right)$ w.r.t. $\tau$. When $n=1$, this iteration is equivalent to DDIM [9]:
$\boldsymbol{f}(\tilde{\boldsymbol{x}}_{t_{i-1}})=\boldsymbol{f}(\tilde{\boldsymbol{x}}_{t_{i}})+h_{t_{i}}\boldsymbol{d}_{\theta}\left(\tilde{\boldsymbol{x}}_{t_{i}},t_{i}\right).$ (6)
where $\tilde{\boldsymbol{x}}_{t}$ depends on the type of $\boldsymbol{d}_{\theta}$. For gradient-based inference, a widely used technique is the finite difference (FD) method [23] as follows: $\boldsymbol{d}_{\theta}^{(k+1)}\left(\boldsymbol{x}_{t},t\right)=\frac{1}{\hat{h}_{t}}\left(\boldsymbol{d}_{\theta}^{(k)}\left(\boldsymbol{x}_{l},l\right)-\boldsymbol{d}_{\theta}^{(k)}\left(\boldsymbol{x}_{t},t\right)\right)+\mathcal{O}(\hat{h}_{t})$. Then, a gradient-based diffusion inference can be derived by using the FD method as follows:
$\vskip-2.84544pt\boldsymbol{f}(\tilde{\boldsymbol{x}}_{t_{i-1}})=\boldsymbol{f}(\tilde{\boldsymbol{x}}_{t_{i}})+h_{t_{i}}\boldsymbol{d}_{\theta}\left(\tilde{\boldsymbol{x}}_{t_{i}},t_{i}\right)+\frac{1}{2}h_{t_{i}}^{2}F_{\theta}(s_{i},{t_{i}}),\vskip-2.84544pt$ (7)
where $F_{\theta}(s_{i},{t_{i}}):=\left(\boldsymbol{d}_{\theta}\left(\tilde{\boldsymbol{x}}_{s_{i}},s_{i}\right)-\boldsymbol{d}_{\theta}\left(\tilde{\boldsymbol{x}}_{t_{i}},t_{i}\right)\right)/\hat{h}_{t_{i}}$ denotes the forward FD, $\hat{h}_{t_{i}}:=\kappa(s_{i})-\kappa(t_{i})$.
DDIM DPM-Solver UniPC DPM-Solver-v3 EVODiff (Ours)
Gradient-based
Bias term (need $\tilde{\boldsymbol{x}}_{0}$)
Variance term
Entropy-aware
Table 1: Strategies employed for optimizing reconstruction error in different methods.
3 Conditional Entropy Reduction in Diffusion Inference
Conditional Entropy in Information Transfer. During DM inference, each iteration reduces uncertainty in intermediate representations via structured denoising. From an information-theoretic view, the information gain between successive states is quantified by the mutual information [24]:
$I_{p}(\boldsymbol{x}_{t_{i}};\boldsymbol{x}_{t_{i+1}})=\displaystyle H_{p}(\boldsymbol{x}_{t_{i}})-\displaystyle H_{p}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}}),$ (8)
where $H_{p}(\boldsymbol{x}_{t_{i}})$ is the entropy of state $\boldsymbol{x}_{t_{i}}$ and $H_{p}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})$ is the conditional entropy of $\boldsymbol{x}_{t_{i}}$ given $\boldsymbol{x}_{t_{i+1}}$. A lower $H_{p}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})$ results in a higher $I_{p}(\boldsymbol{x}_{t_{i}};\boldsymbol{x}_{t_{i+1}})$, which suggests that a well-designed method effectively utilizes the information from $\boldsymbol{x}_{t_{i+1}}$ to refine the estimate of $\boldsymbol{x}_{t_{i}}$.
Conditional Variance and Conditional Entropy. We denote $p(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}},\boldsymbol{x}_{0})$ as $p(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})$, and $\mathrm{Var}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})$ as the conditional variance $\mathrm{Var}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}},\boldsymbol{x}_{0})$ for brevity. In DMs [2, 3], the reverse transition $p(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}},\boldsymbol{x}_{0})$ is commonly approximated as a Gaussian distribution under the Markov assumption, i.e., $p(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})\approx\mathcal{N}(\boldsymbol{\mu}_{t_{i}},\boldsymbol{\Sigma}_{t_{i}})$, simplifying both model training and theoretical analysis [21, 25, 26, 12]. Accordingly, the conditional entropy $H_{p}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})$ simplifies to:
$H_{p}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})\approx C+1/2\cdot\log\det(\mathrm{Var}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})),$ (9)
where $C=1/2\cdot d(\log 2\pi+1)$. Thus, $H_{p}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})$ is intrinsically tied to its conditional variance:
$H_{p}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})\propto\log\det(\mathrm{Var}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})).$ (10)
This indicates that minimizing conditional variance directly reduces conditional entropy.
Reconstruction Error and Conditional Variance. In the forward process, $q(\boldsymbol{x}_{t})$ approaches a standard Gaussian as $t$ increases, but $q(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})$ remains structured around scaled versions of $\boldsymbol{x}_{0}$. Since $H_{q}(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})\leq H_{q}(\boldsymbol{x}_{t})$, information about $\boldsymbol{x}_{0}$ persists in the modeled $\boldsymbol{x}_{t}$, and the inference process seeks to recover it. Let $\boldsymbol{\mu}_{t_{i}|t_{i+1}}=\mathbb{E}_{q}[\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}}]$. The MSE between the inference states and its posterior mean is $\mathbb{E}_{q}[\|\boldsymbol{x}_{t_{i}}-\boldsymbol{\mu}_{t_{i}|t_{i+1}}\|^{2}]=\text{Tr}(\mathrm{Var}_{q}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}}))$. Leveraging the connection with conditional variance, under the isotropic assumption commonly used in DMs, we obtain:
$\min H_{q}(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})\Leftrightarrow\min\mathbb{E}_{q}[\|\boldsymbol{x}_{t_{i}}-\boldsymbol{\mu}_{t_{i}|t_{i+1}}\|^{2}].$ (11)
We now decompose the reconstruction error between $\boldsymbol{x}_{t_{i}}$ and $\boldsymbol{x}_{0}$ using the law of total expectation.
Proposition 3.1.
Note that $\boldsymbol{x}_{t_{i}}-\boldsymbol{x}_{0}=(\boldsymbol{x}_{t_{i}}-\boldsymbol{\mu}_{t_{i}|t_{i+1}})+(\boldsymbol{\mu}_{t_{i}|t_{i+1}}-\boldsymbol{x}_{0})$, we have
$\mathbb{E}_{q}\left[\|\boldsymbol{x}_{t_{i}}-\boldsymbol{x}_{0}\|^{2}\right]=\underbrace{\mathbb{E}_{q}\left[\|\boldsymbol{x}_{t_{i}}-\boldsymbol{\mu}_{t_{i}|t_{i+1}}\|^{2}\right]}_{\textbf{Variance term}}+\underbrace{\mathbb{E}_{q}\left[\|\boldsymbol{\mu}_{t_{i}|t_{i+1}}-\boldsymbol{x}_{0}\|^{2}\right]}_{\textbf{Bias term}}.$ (12)
where details of this reconstruction error decomposition are provided in Appendix LABEL:proofreconerror.
This decomposition reveals two distinct inference approaches for reducing the reconstruction error: (1) minimizing the conditional variance term directly; and (2) optimizing the bias error term with the prior $\boldsymbol{x}_{0}$, which is often approximated by deterministic DM samplers. Although the total reconstruction error includes both the variance and bias terms, optimizing conditional variance becomes the main actionable mechanism, since we do not often have access to the desired $\boldsymbol{x}_{0}$ during inference. Thus, optimizing conditional variance is central to inference; various methods are summarized in Table 1.
(a) Quantitative results of FID $\downarrow$ show that efficient entropy reduction (RE) method consistently improves image quality compared to FD-based method in Eq.(7) across various ablation scenarios.
Variance-Driven Conditional Entropy Analysis. We demonstrate how entropy reduction effectively steers samples toward the target distribution, supported by both theory and empirical evidence. Theoretically, DM denoising functions as an entropy reduction mechanism grounded in Langevin dynamics [19] and non-equilibrium thermodynamics [18]. This principle dictates that more efficient entropy reduction will accelerate convergence by steering samples toward high-probability regions of the target distribution. Figure 1 visually illustrates this, showing the trajectories of DDIM versus our gradient-based inference. Further empirical evidence is presented in Figure 3(a), which details an ablation study on CIFAR-10 and CelebA-64 comparing our entropy reduction-focused (RE) method with traditional FD-based gradients.
Our analysis reveals that gradient-based methods excel at driving entropy reduction by optimizing conditional variance, which efficiently guides noisy states toward the desired distribution. For theoretical tractability, we assume independence between estimated noise at different timesteps, in line with DDPM’s training objective of independent MSE minimization. While neural network parameter-sharing during training could introduce dependencies, prior works like [2, 9] justify this surrogate by showing that these dependencies have a negligible performance impact.
We identify two sources of conditional entropy in the reverse transition of DMs: the inference path uncertainty (e.g., from ODE/SDE solvers) and the model’s own intrinsic uncertainty. While simple ODE-solver paths address the former, they may not effectively reduce the entropy contributed by the model. We therefore propose an approach that moves beyond these simple paths to target the total conditional entropy. From this perspective, we first derive Proposition 3.2 (Proof in Appendix LABEL:prorelativere).
Proposition 3.2.
The gradient-based inference in Eq. (7) can reduce conditional entropy more efficiently than the first-order inference in Eq. (6) when $\frac{h_{t_{i}}}{\hat{h}_{t_{i}}}\in\left[1,\frac{4\cdot\mathrm{Var}(\boldsymbol{\epsilon}_{\theta}(\tilde{\boldsymbol{x}}_{t_{i}},t_{i})}{\mathrm{Var}(\boldsymbol{\epsilon}_{\theta}(\tilde{\boldsymbol{x}}_{s_{i}},s_{i})+\mathrm{Var}(\boldsymbol{\epsilon}_{\theta}(\tilde{\boldsymbol{x}}_{t_{i}},t_{i})}\right]$.
This shows that gradient-based inference can achieve larger reductions in uncertainty when the step size ratio is appropriately chosen. A practical interval for Proposition 3.2 is discussed in Remark LABEL:modiferi.
Furthermore, we identify that solvers like DPM-Solver [13] and the Heun iterations in EDM [12] can be understood through conditional entropy reduction, with details in the Appendix LABEL:erasomesolver.
Proposition 3.3.
The acceleration mechanisms of DPM-Solver and the Heun iterations in EDM can be unified and explained as specific implementations of the conditional entropy reduction framework. Finally, we theoretically establish why denoising iterations using data prediction perform better than those using noise prediction. The proof is provided in Appendix LABEL:profdatanoise.
Theorem 3.4.
Data prediction parameterization reduces reconstruction errors more effectively than its noise counterpart. Under independence assumptions, it also reduces conditional entropy.
In summary, we provide a variance-driven conditional entropy analysis for diffusion inference, which theoretically explains the superior performance of gradient-based inference and data prediction parameterization. In particular, Theorem 3.4 demonstrates that data parameterization, by directly targeting the data distribution, avoids the error-prone chain of $\boldsymbol{\epsilon}_{t}\mapsto\boldsymbol{x}_{t}\mapsto\boldsymbol{x}_{0}$.
Algorithm 1 EVODiff: Optimizing Denoising Variance of Diffusion Model Inference.
0: initial $\boldsymbol{x}_{T}$, time schedule $\left\{t_{i}\right\}_{i=0}^{N}$, model $\boldsymbol{x}_{\theta}$.
[…]
2: Denote $\boldsymbol{g}(\boldsymbol{x}_{t_{i}}):=\frac{\sigma_{t_{i-1}}}{\sigma_{t_{i}}}\boldsymbol{x}_{t_{i}}+\sigma_{t_{i-1}}h_{t_{i}}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)$. # Euler’s or DDIM’s iteration.
3:for $i\leftarrow$ NN to 11 do
4:  $\boldsymbol{x}_{t_{i}}\leftarrow\boldsymbol{g}(\boldsymbol{x}_{t_{i+1}})$.
5:  $\boldsymbol{x}_{t_{i-1}}\leftarrow\boldsymbol{g}(\boldsymbol{x}_{t_{i}})+\sigma_{t_{i-1}}\frac{h_{t_{i}}^{2}}{2}B_{\theta}(t_{i},l_{i})$.
6:  $B_{\theta}({t_{i}})\leftarrow\left(1-\frac{\eta_{i}}{2}\right)B_{\theta}(s_{i},t_{i})+\frac{\eta_{i}}{2}B_{\theta}(t_{i},l_{i})$, where $\eta_{i}$ is refined by Eq. (25).
7:  $\boldsymbol{x}_{t_{i-1}}\leftarrow\boldsymbol{g}(\boldsymbol{x}_{t_{i}})+\sigma_{t_{i-1}}\frac{h_{t_{i}}^{2}}{2\zeta_{i}}B_{\theta}({t_{i}})$, where $\zeta_{i}$ is refined by Eq. (25).
8:end for
[…]
4 Optimizing Diffusion Model Inference via Entropy-aware Variance Control
4.1 Denoising Variance Analysis for Gradient-based Inference
Our focus is on multi-step iterations using data parameterization, which has shown superiority through the theoretical result in Theorem 3.4 and prior empirical evidence from [13, 20].
Note that $\boldsymbol{f}(\boldsymbol{x}_{t})=\frac{\boldsymbol{x}_{t}}{\sigma_{t}}$ for the ODE for data prediction defined in Eq. (5), $\boldsymbol{\iota}(\boldsymbol{x}_{t_{i-1}})=\frac{\boldsymbol{x}_{t_{i-1}}}{\sigma_{t_{i-1}}}-\frac{\boldsymbol{x}_{t_{i}}}{\sigma_{t_{i}}}$. Formally, the multi-step iteration can be written as:
$\frac{\boldsymbol{x}_{t_{i-1}}}{\sigma_{t_{i-1}}}-\frac{\boldsymbol{x}_{t_{i}}}{\sigma_{t_{i}}}=h_{t_{i}}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)+\frac{1}{2}h_{t_{i}}^{2}B_{\theta}(t_{i},{t_{i+1}}),$ (13)
where $B_{\theta}({t_{i}},t_{i+1}):=(\boldsymbol{x}_{\theta}(\boldsymbol{x}_{t_{i}},t_{i})-\boldsymbol{x}_{\theta}(\boldsymbol{x}_{t_{i+1}},t_{i+1}))/h_{t_{i+1}}$ denotes the backward FD. In this iteration, the smaller step size $|h_{t_{i}}|$ compared to $|h_{t_{i}}-h_{t_{i+1}}|$ in the single-step case (Appendix LABEL:singleanaly) reduces the conditional entropy, offering greater potential to improve the denoising process.
Denote $\bar{\zeta}_{i}=(1-\zeta_{i})$. A straightforward improvement for Eq. (13) can be formulated as follows:
$\frac{\boldsymbol{x}_{t_{i-1}}}{\sigma_{t_{i-1}}}-\frac{\boldsymbol{x}_{t_{i}}}{\sigma_{t_{i}}}=h_{t_{i}}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)+\frac{1}{2}h_{t_{i}}^{2}B_{\theta}(t_{i},l_{i}),$ (14)
where $\boldsymbol{x}_{\theta}(\boldsymbol{x}_{l_{i}},l_{i}):=\zeta_{i}\boldsymbol{x}_{\theta}(\boldsymbol{x}_{t_{i}},t_{i})+\bar{\zeta}_{i}\boldsymbol{x}_{\theta}(\boldsymbol{x}_{t_{i+1}},t_{i+1})$ represents a linear interpolation. Similarly, the implicit approach is: $\frac{\boldsymbol{x}_{t_{i-1}}}{\sigma_{t_{i-1}}}-\frac{\boldsymbol{x}_{t_{i}}}{\sigma_{t_{i}}}=h_{t_{i}}\boldsymbol{x}_{\theta}(\boldsymbol{x}_{t_{i}},t_{i})+\frac{1}{2}h_{t_{i}}^{2}B_{\theta}(s_{i},t_{i})$, where $\boldsymbol{x}_{\theta}(\boldsymbol{x}_{s_{i}},s_{i}):=\zeta_{i}\boldsymbol{x}_{\theta}(\boldsymbol{x}_{t_{i-1}},t_{i-1})+\bar{\zeta}_{i}\boldsymbol{x}_{\theta}(\boldsymbol{x}_{t_{i}},t_{i})$. Note that these two improvement approaches can be unified as
$\frac{\boldsymbol{x}_{t_{i-1}}}{\sigma_{t_{i-1}}}-\frac{\boldsymbol{x}_{t_{i}}}{\sigma_{t_{i}}}=h_{t_{i}}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)+\frac{1}{2}h_{t_{i}}^{2}\zeta_{i}\bar{B}_{\theta}(t_{i};u_{i}),$ (15)
where $\bar{B}_{\theta}(t_{i};u_{i})=B_{\theta}(s_{i},t_{i})$ when $u_{i}=s_{i}$, and $\bar{B}_{\theta}(t_{i};u_{i})=B_{\theta}(t_{i},l_{i})$ when $u_{i}=l_{i}$.
Remark 4.1.
The RE-based multi-step iterations in Eq. (15) reduce the conditional entropy of iteration in Eq. (13) by leveraging model parameters from low-variance regions.
We provide the convergence guarantees for the RE-based multi-step iteration described in Eq. (15).
Theorem 4.2.
If $\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t},t\right)$ satisfies Assumption LABEL:assumption, the RE-based multi-step iteration constitutes a globally convergent second-order iterative algorithm. The proof is provided in Appendix LABEL:mutidataconver.
However, a key question arises: how should $\zeta_{i}$ and $\hat{h}_{t_{i}}$ be determined? As $\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)$ is predict the clean data $\boldsymbol{x}_{0}$ from the noisy data $\boldsymbol{x}_{t_{i}}$, we present a remark for ideal cases.
Remark 4.3.
If $\mathrm{Var}\left(\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)\right)\propto\sigma_{t_{i}}^{2}$, the variance minimization can be achieved by setting $\zeta_{i}=\sigma_{t_{i-1}}^{2}/(\sigma_{t_{i}}^{2}+\sigma_{t_{i-1}}^{2})$ for $\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{s_{i}},s_{i}\right)$ and $\zeta_{i}=\sigma_{t_{i}}^{2}/(\sigma_{t_{i}}^{2}+\sigma_{t_{i+1}}^{2})$ for $\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{l_{i}},l_{i}\right)$.
Furthermore, we can improve the iteration of Eq. (13) by incorporating $B_{\theta}(t_{i},s_{i})$ and $B_{\theta}(s_{i},t_{i})$:
$\frac{\boldsymbol{x}_{t_{i-1}}}{\sigma_{t_{i-1}}}-\frac{\boldsymbol{x}_{t_{i}}}{\sigma_{t_{i}}}=h_{t_{i}}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)+\frac{1}{2}h_{t_{i}}^{2}\left((1-\eta_{i})B_{\theta}(s_{i},t_{i})+\eta_{i}B_{\theta}(t_{i},l_{i})\right),$ (16)
where $\eta_{i}$ determines the gradient term variance. Therefore, from the lens of conditional entropy reduction, we can establish an optimization objective to directly minimize this variance.
            NFE=55            NFE=88            NFE=$10$            NFE=$12$
DPM-Solver ++ [20] Refer to caption Refer to caption Refer to caption Refer to caption
EVODiff1 Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Sample comparison of our method vs. baseline using the pre-trained EDM on CIFAR-10.
4.2 Optimizing Denoising Variance with Evolution State Differences
We observe that the conditional variance in gradient-based iterations can be composed of two critical components: the variance of the gradient estimation term itself and the variance between the gradient term and the first-order term. Specifically, the unified iteration in Eq. (15) can be rewritten as:
$\frac{\boldsymbol{x}_{t_{i-1}}}{\sigma_{t_{i-1}}}-\frac{\boldsymbol{x}_{t_{i}}}{\sigma_{t_{i}}}=h_{t_{i}}\left(\left(1\mp\frac{\zeta_{i}}{2}\frac{h_{t_{i}}}{h_{\mu_{i}}}\right)\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)\pm\frac{\zeta_{i}}{2}\frac{h_{t_{i}}}{h_{\mu_{i}}}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{\mu_{i}},\mu_{i}\right)\right).$ (17)
As $\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{\mu_{i}},\mu_{i}\right)$ and $\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)$ are known in multi-step iterative mechanisms, the $\mathrm{Var}(\boldsymbol{x}_{t_{i-1}}|\boldsymbol{x}_{t_{i}})$ is controlled by the value of $\zeta_{i}$. Inspired by Eq. (17), we observe that $\zeta_{i}$ balances the variance between the gradient term and the first-order term. By harmonizing their statistical characteristics, we can efficiently reduce the conditional variance. To achieve this, we define $G(\zeta_{i})=\left(1-\zeta_{i}\right)\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)+\zeta_{i}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{\mu_{i}},\mu_{i}\right)$ to balance the variance between two terms. Moreover, the variance of the gradient term itself should be balanced by $\eta_{i}$, as we should optimize $\eta_{i}$ in Eq. (16).
(a) FID $\downarrow$ scores for gradient-based inference methods on ImageNet-64 and FFHQ-64.
Refining $\zeta_{i}$ with Evolution State Differences. Our goal is to refine $\zeta_{i}$ using the available information at current step. As discussed above, we can optimize $\zeta_{i}$ to control the conditional variance by formulating an objective involving $G(\zeta_{i})$. On the one hand, we can rewrite the iteration in Eq. (17) as
$\hat{\boldsymbol{x}}_{1,t_{i}}=\frac{\sigma_{t_{i}}}{\sigma_{t_{i-1}}}\boldsymbol{x}_{t_{i-1}}-\sigma_{t_{i}}h_{t_{i}}G(\zeta_{i}).$ (18)
Notice that $\tilde{\boldsymbol{x}}_{t_{i}}$ in Eq. (18) is determined by $\zeta_{i}$. On the other hand, we can consider $\boldsymbol{x}_{t_{i-1}}$ as a starting point and perform an inverse iterative from $t_{i-1}$ to $t_{i}$ to approximate $\boldsymbol{x}_{t_{i}}$ as follows:
$\frac{\boldsymbol{x}_{t_{i}}}{\sigma_{t_{i}}}-\frac{\boldsymbol{x}_{t_{i-1}}}{\sigma_{t_{i-1}}}=\int_{\kappa(t_{i-1})}^{\kappa(t_{i})}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{\psi(\tau)},\psi(\tau)\right)\mathrm{d}\tau.$ (19)
Similar to the Eq. (13), this inverse estimation of Eq. (19) is as follow:
$\hat{\boldsymbol{x}}_{2,t_{i}}=\frac{\sigma_{t_{i}}}{\sigma_{t_{i-1}}}\boldsymbol{x}_{t_{i-1}}-\sigma_{t_{i}}h_{t_{i}}\boldsymbol{x}_{\theta}\left(\boldsymbol{x}_{t_{i-1}},t_{i-1}\right)+\sigma_{t_{i}}\frac{1}{2}h_{t_{i}}^{2}B_{\theta}(s_{i},{t_{i}}).$ (20)
Drawing from equations (18) and (20), we can determine $\zeta_{i}$ by minimizing the differences between two estimations. Then, the optimization objective for $\zeta_{i}$ is defined as follows:
$\min\limits_{\zeta_{i}>0}~\mathcal{L}_{1}(\zeta_{i}):=\left\|(\hat{\boldsymbol{x}}_{1,t_{i}}-\boldsymbol{x}_{t_{i}})+(\hat{\boldsymbol{x}}_{2,t_{i}}-\boldsymbol{x}_{t_{i}})\right\|,$ (21)
Directly solving this objective is challenging, as the optimal $\boldsymbol{x}_{t_{i}}$ is unknown. Fortunately, as $\mathcal{L}_{1}(\zeta_{i})\leq\|\hat{\boldsymbol{x}}_{1,t_{i}}+\hat{\boldsymbol{x}}_{2,t_{i}}\|+\left\|2\boldsymbol{x}_{t_{i}}\right\|$, we observe that $\|2\boldsymbol{x}_{t_{i}}\|$ is independent of the $\zeta_{i}$. Then, we can use $\tilde{\boldsymbol{x}}_{t_{i}}$ to replace $\boldsymbol{x}_{t_{i}}$ as when optimizing $\mathcal{L}_{1}(\zeta_{i})$. Denote $P(\boldsymbol{x}_{t_{i-1}}):=\hat{\boldsymbol{x}}_{2,t_{i}}+\frac{\sigma_{t_{i}}}{\sigma_{t_{i-1}}}\boldsymbol{x}_{t_{i-1}}-2\tilde{\boldsymbol{x}}_{t_{i}}$ for brevity. Then, $\mathcal{L}_{1}(\zeta_{i})$ can be rewritten as: $\mathcal{L}_{1}(\zeta_{i})=\|P(\boldsymbol{x}_{t_{i-1}})-\sigma_{t_{i}}h_{t_{i}}G(\zeta_{i})\|$.
Lemma 4.4.
When the constraint on $\zeta_{i}$ is relaxed, $\min\limits_{\zeta_{i}}\mathcal{L}^{2}_{1}(\zeta_{i})$ possesses the closed-form solution: $\zeta^{\ast}_{i}=-(\text{vec}^{T}(D_{i})\text{vec}(\tilde{P}_{i}))\big/(\sigma_{t_{i}}h_{t_{i}}\text{vec}^{T}(D_{i})\text{vec}(D_{i})),$ (22) where $\tilde{P}_{i}:=P(\boldsymbol{x}_{t_{i-1}})-\sigma_{t_{i}}h_{t_{i}}\boldsymbol{x}_{\theta}\left(\tilde{\boldsymbol{x}}_{t_{i}},t_{i}\right)$, $D_{i}:=\boldsymbol{x}_{\theta}\left(\tilde{\boldsymbol{x}}_{t_{i-1}},t_{i-1}\right)-\boldsymbol{x}_{\theta}\left(\tilde{\boldsymbol{x}}_{t_{i}},t_{i}\right)$, and $\text{vec}(\cdot)$ denotes the vectorization operation. The proof is provided in Appendix LABEL:lemmaproof.
Refining $\eta_{i}$ by Balancing Gradient Errors. Our goal is to refine $\eta_{i}$ using the available information at current step. Denote $\tilde{\Delta}^{g}_{t_{i}}=(1-\eta_{i})B_{\theta}(s_{i},t_{i})+\eta_{i}B_{\theta}(t_{i},l_{i})$, $E(t_{i-1},t_{i}):=B_{\theta}(s_{i},t_{i})-\boldsymbol{x}_{\theta}^{(1)}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)$ as gradient error. For balancing this errors, we formulate the following optimization objective:
$\min\limits_{\eta_{i}\in(0,1]}\mathcal{L}_{2}(\eta_{i}):=\left\|(1-\eta_{i})E(t_{i-1},t_{i})+\eta_{i}E(t_{i},t_{i+1})\right\|.$ (23)
We can rewrite $\mathcal{L}_{2}(\eta_{i})$ as $\mathcal{L}_{2}(\eta_{i})=\|\tilde{\Delta}^{g}_{t_{i}}-\boldsymbol{x}_{\theta}^{(1)}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)\|$. Denote $\mathcal{L}_{2\text{s}}(\eta_{i}):=\|\tilde{\Delta}^{g}_{t_{i}}\|$. Then, $\mathcal{L}_{2}(\eta_{i})\leq\mathcal{L}_{2\text{s}}(\eta_{i})+\|\boldsymbol{x}_{\theta}^{(1)}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)\|$, where $\boldsymbol{x}_{\theta}^{(1)}\left(\boldsymbol{x}_{t_{i}},t_{i}\right)$ can be regarded as a specific constant term independent of the target $\eta_{i}$. Thus, the $\eta_{i}$ can be obtained by optimizing the tractable $\mathcal{L}_{2\text{s}}(\eta_{i})$.
Lemma 4.5.
When the constraint on $\eta_{i}$ is relaxed, $\min\limits_{\eta_{i}}\mathcal{L}^{2}_{2\text{s}}(\eta_{i})$ possesses the closed-form solution: $\eta^{\ast}_{i}=-(\text{vec}^{T}(\tilde{B}_{i})\text{vec}(B_{\theta}(t_{i},l_{i})))\big/(\text{vec}^{T}(\tilde{B}_{i})\text{vec}(\tilde{B}_{i})),$ (24) where $\tilde{B}_{i}:=B_{\theta}(s_{i},t_{i})-B_{\theta}(t_{i},l_{i})$. The proof is similar to that of Lemma 4.4.
From Lemmas 4.4 and 4.5, $\mathcal{L}_{1\text{s}}(\zeta_{i})$ and $\mathcal{L}_{2\text{s}}(\eta_{i})$ have closed-form solutions when the constraints are relaxed. Then, the parameters in Algorithm 1 are derived by mapping these solutions as follows:
$\eta_{i}=\text{Sigmoid}\left(|\eta_{i}^{\ast}|\right),~\zeta_{i}=\text{Sigmoid}\left(-(|\zeta_{i}^{\ast}|-\mu)\right),$ (25)
where $\zeta_{i}^{\ast}$ and $\eta_{i}^{\ast}$ are defined in Eqs. (22) and (24), respectively, and $\mu$ is the shift parameter. This optimization-driven approach captures the state differences during iteration while avoiding the computational cost of constrained optimization problems [27]. Additional details are provided in Appendix LABEL:addmultiiter. Ablation studies on $\mu$ are provided in Table LABEL:tab:ablation_shift_parameters of Appendix LABEL:AblaEvoDiff.
Finally, we prove the global convergence of the proposed inference method for data parameterization.
Theorem 4.6.
The diffusion inference method in Algorithm 1 exhibits second-order global convergence with a local error of $\mathcal{O}(h_{t_{i}}^{3})$. The proof is provided in Appendix LABEL:converesampling.
5 Experiments
We experimentally validate our method on a diverse suite of DMs and datasets, including CIFAR-10, CelebA-64, FFHQ-64, ImageNet-64, ImageNet-256, and LSUN-Bedrooms. Our evaluation uses standard metrics such as Fréchet Inception Distance (FID) and Inception Score (IS) across a varying number of function evaluations (NFEs). We also compare our method on Stable Diffusion v1.4 and v1.5 using CLIP and Aesthetic scores. Our evaluation focuses on the data prediction parameterization on different diffusion-based generative models including pixel-space diffusion and latent-space diffusion. Additional experimental details and results are provided in Appendix LABEL:appexp.
Table 2: Quantitative results of FID $\downarrow$ and IS $\uparrow$ scores for gradient-based methods on ImageNet-256, FFHQ-64, and CIFAR-10. The results are evaluated on 10k and 50k samples for various NFEs. The DPM-Solver++ is our baseline. Error optimization strategies across methods are shown in Table 1.
Model Method/NFE Entropy-aware? 5 8 10 12
[…]
EVODiff (our) 17.84 7.89 3.98 9.37 2.78 9.64 2.30 9.80
Model Method/NFE Entropy-aware? 5 10 15 20
[…]
EVODiff (our) $\checkmark$ 19.65 3.18 5.31 3.32 3.04 3.35 2.66 3.38
Model Method/NFE Reference-based? 5 10 15 20
[…]
DPM-Solver-v3 $\checkmark$ 14.92 105.85 8.14 146.82 7.70 153.79 7.42 154.35
[…]
NFE=25 NFE=35
DPM-Solver ++ [20] Refer to caption Refer to caption
UniPC [16] Refer to caption Refer to caption
EVODiff 1 Refer to caption Refer to caption
Figure 6: Random samples from the Stable-Diffusion-v1.5 model [28] with a guidance scale of 7.5, using varying NFEs and the prompt “Giant caterpillar riding a bicycle". Even at a low 25 NFE, EVODiff produces high-fidelity, semantically correct images while competing methods fail with severe artifacts, demonstrating the superiority of our entropy-aware variance optimized method.
Table 3: Ablation study of our variance-driven gradient-based approach applied to DMs with data parameterization on ImageNet-256 and CIFAR-10. The DPM-Solver++ is our baseline method.
Method Model NFE
55 66 88 $10$ $12$ $15$ 2020
Baseline CIFAR-10 $27.96$ $16.87$ $8.40$ $5.10$ $3.70$ $2.83$ $2.33$
RE-based in Eq. (15) $21.39$ $12.14$ $4.81$ $2.98$ $2.44$ $2.15$ $2.08$
EVODiff 17.83 9.17 3.98 2.78 2.30 2.12 2.06
Baseline ImageNet-256 $16.62$ $12.86$ $9.73$ $8.68$ $8.17$ $7.80$ $7.51$
Eq. (15) with balanced $\zeta_{i}$ $15.31$ $12.14$ $9.46$ $8.57$ $8.07$ $7.76$ $7.48$
Eq. (16) with refined $\zeta_{i}$ 13.80 10.91 $8.91$ $8.23$ $7.89$ $7.58$ $7.36$
EVODiff ($r_{\text{logSNR}}$) $13.98$ $10.98$ $\mathbf{8.84}$ $8.16$ $7.81$ $7.52$ $7.32$
EVODiff ($r_{\text{refined}}$) $14.33$ $11.16$ $8.95$ $\mathbf{8.14}$ $\mathbf{7.79}$ $\mathbf{7.48}$ $\mathbf{7.25}$
Table 4: FID score ($\downarrow$) and generation time comparison between EvoDiff, DPM-Solver++ (DPM++ for short), and UniPC. All methods were evaluated on a latent diffusion model [28] trained on the LSUN-Bedrooms dataset with 50k samples.
FID Score ($\downarrow$) Generation Time (s)
NFE DPM++(2m) DPM++(3m) UniPC(3m) EvoDiff Gain DPM++ EvoDiff Gain
5 21.286 18.611 13.969 7.912 43.4% 3577.6 3488.4 -89.2 (2.5%)
6 10.966 8.519 6.556 4.909 25.1% 3800.6 3719.4 -81.2 (2.1%)
8 5.127 4.148 3.963 3.756 5.2% 4273.3 4046.9 -226.4 (5.3%)
10 3.881 3.607 3.563 3.332 6.5% 4746.7 4699.6 -47.1 (1.0%)
12 3.516 3.429 3.357 3.084 8.1% 4703.8 4678.1 -25.7 (0.5%)
15 3.341 3.284 3.182 2.918 8.3% 5973.1 5913.5 -59.6 (1.0%)
20 3.251 3.167 3.075 2.853 7.2% 7238.4 7154.2 -84.2 (1.2%)
Table 5: Single-batch sample quality comparison for text-to-image generation, measured by CLIP score ($\uparrow$) and Aesthetic score ($\uparrow$) using Stable Diffusion v1.4 and v1.5 under identical settings.
model method NFN=10 NFN=25
CLIP Aesthetic CLIP Aesthetic
Average maximum Average maximum Average maximum Average maximum
sd-v1.4 DPM-Solver++ 33.07 34.12 5.71 5.83 32.66 36.12 5.74 5.92
EVODiff 33.79 35.00 5.77 5.87 32.83 34.50 5.79 5.97
sd-v1.5 DPM-Solver++ 32.98 35.84 5.74 5.99 32.54 34.70 5.79 5.92
EVODiff 33.07 36.62 5.75 5.98 32.53 34.72 5.81 5.99
To validate the contributions of each component of EVODiff, we performed a constructive ablation study (Table 3), which builds upon our initial finding that entropy-reduction (RE) based methods consistently outperform traditional FD-based approaches (Figure 3(a)). Starting from a baseline second-order solver, we incrementally introduced our entropy-reduction (RE-based) formulation and the final evolution-state-driven parameter optimization. Each step demonstrates a clear improvement in FID, culminating in the full EVODiff algorithm which consistently achieves the best performance. Further detailed ablations on hyperparameters such as the step-size ratio $r_{i}$ and the shift parameter $\mu$ can be found in Appendix LABEL:AblaEvoDiff (Tables LABEL:ablation256withbaseline and LABEL:tab:ablation_shift_parameters in of Appendix LABEL:AblaEvoDiff), confirming the robustness of our method. Additionally, we demonstrate the generalizability of our core principles by applying them to enhance other frameworks like DPM-Solver-v3, with results shown in Table LABEL:tab:recifar10 and Figure LABEL:fig:v3cifar105nfe.
We evaluate our method against advanced gradient-based solvers, including DPM-Solver++ [20], DEIS [14], UniPC [16], and DPM-Solver-v3 [17] on CIFAR-10, FFHQ-64, ImageNet-64, and ImageNet-256 datasets. The results in Tables 2 and LABEL:ablation256withbaseline and Figure 6(a) consistently demonstrate the superior performance of EVODiff. Furthermore, we evaluate our method with available logSNR and EDM noise schedules, further validating its consistently robust performance. The results are shown in Tables LABEL:tab:logsnr_cifar10, LABEL:tab:imageNet64_logSNR, LABEL:tab:imageNet64_edm, LABEL:tab:edm_ffhq, LABEL:tab:logsnr_ffhq. We also evaluate our method on the text-to-image generation task, as shown in Table 5. Figures 4 and 6 provide a visual comparison of a generated sample. Finally, Table LABEL:tab:computation_comparison compares the inference time at various NFEs on ImageNet-256 between our method and the baseline.
Finally, beyond pixel-space DMs, we evaluate EVODiff against advanced gradient-based solvers on popular latent-space DMs. On the LSUN-Bedrooms dataset (Table 4), EVODiff consistently achieves the best FID scores across all NFE settings, with particularly significant improvements at low NFE counts (43.4% reduction at 5 NFE, from 13.969 to 7.912 compared to UniPC). Notably, EVODiff also reduces generation time by up to 5.3% while maintaining superior quality. For text-to-image generation using Stable Diffusion v1.4 and v1.5 (Table 5), EVODiff achieves competitive CLIP scores and the best Aesthetic scores compared to the strong DPM-Solver++ baseline, demonstrating its effectiveness in preserving both semantic alignment and visual quality.
Conclusions
In this work, we propose EVODiff, a novel inference-time refinement method based on entropy-aware variance optimization. It significantly improves both efficiency and generative quality without relying on reference trajectories. Specifically, our work first establishes a principled, information-theoretic foundation that explains why data-prediction parameterization outperforms its noisy counterpart and demonstrates how optimizing conditional variance reduces transition and reconstruction errors without relying on reference trajectories. Building on these insights, EVODiff systematically reduces uncertainty in each denoising step, thereby accelerating convergence and significantly improving sample quality. Extensive experiments demonstrate EVODiff’s effectiveness across diverse settings with SOTA performance: on CIFAR-10, it outperforms the DPM-Solver++ baseline by 45.5% at 10 NFE (from 5.10 to 2.78 FID); on ImageNet-256, it reduces NFE cost by 25% (from 20 to 15 NFE) while maintaining high-fidelity generation; on LSUN-Bedrooms, it achieves up to 43.4% FID improvement over UniPC with 5.3% faster generation; and for text-to-image generation with Stable-Diffusion models, it produces superior visual quality while preserving semantic alignment. Our method achieves SOTA results and establishes a reference-free, variance-controlled inference framework, effectively addressing the trade-off between sampling efficiency and generative quality.
Limitations and Broader Impacts
A limitation of EVODiff is that it currently relies on the data-prediction parameterization for diffusion model inference. Additionally, leveraging information-theoretic principles to enhance inference efficiency and optimize information flow during the sampling process remains a promising direction for future research. While EVODiff improves both generation quality and efficiency, we acknowledge its dual-use nature, similar to that of other generative models.
Acknowledgments and Disclosure of Funding
This work was supported in part by grants from National Natural Science Foundation of China (52539005), the fundamental research program of Guangdong, China (2023A1515011281), the China Scholarship Council (202306150167), Guangdong Basic and Applied Basic Research Foundation (24202107190000687), Foshan Science and Technology Research Project(2220001018608).
Appendix
Appendix A List of Notations
Table 6: Random samples from Stable-Diffusion-v1.5 with a guidance scale 7.5, using varying NFEs and the prompt “A robot chef cooking a meal in a futuristic kitchen, with glowing utensils and a holographic recipe book, highly detailed, sci-fi atmosphere". Our EVODiff inference improves content continuity and consistency, and effectively reduces visual artifacts.
$t,s,l,T$ Time variables and endpoint in the diffusion process, $t\in[0,T]$.
$i,j,k$ Indices for discrete time steps.
$\lambda(t)$ Log-Signal-to-Noise Ratio (log-SNR) time, defined as $\log(\alpha_{t}/\sigma_{t})$.
$\boldsymbol{x}_{t},\boldsymbol{x}_{t_{i}},\tilde{\boldsymbol{x}}_{t}$ State vector at time $t$ (continuous, discrete, or approximate).
$\boldsymbol{x}_{0},\boldsymbol{x}_{T}$ Endpoints: $\boldsymbol{x}_{0}\sim q(\boldsymbol{x}_{0})$, $\boldsymbol{x}_{T}\sim\mathcal{N}(\boldsymbol{0},\hat{\sigma}^{2}\boldsymbol{I})$.
$\alpha_{t},\sigma_{t}$ Noise schedule parameters (signal preservation $\alpha_{t}$, noise level $\sigma_{t}$).
$\boldsymbol{\epsilon}$ Random noise sampled from $\mathcal{N}(\mathbf{0},\boldsymbol{I})$.
$\boldsymbol{\omega}_{t},\overline{\boldsymbol{\omega}}_{t}$ Forward and reverse Wiener processes.
SNR Signal-to-noise ratio: $\alpha_{t}^{2}/\sigma_{t}^{2}$.
$\theta$ Parameters of the neural network model.
$\boldsymbol{\epsilon}_{\theta}(\cdot),\boldsymbol{x}_{\theta}(\cdot),\boldsymbol{d}_{\theta}(\cdot)$ Network predictions: noise, clean data,  unified noise and data.
$\boldsymbol{d}_{\theta}^{(k)}(\cdot)$ $k$-th order derivative of $\boldsymbol{d}_{\theta}$ w.r.t. $\tau$.
$q(\cdot)$ Forward process distributions, e.g., $q(\boldsymbol{x}_{0})$, $q(\boldsymbol{x}_{t})$, $q(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})$.
$p(\boldsymbol{x}_{t_{i}}|\boldsymbol{x}_{t_{i+1}})$ Reverse transition distribution.
$\mathcal{N}(\mu,\Sigma),\mathcal{U}(a,b)$ Gaussian distribution with mean $\mu$, covariance $\Sigma$; uniform distribution on $[a,b]$.
$f(t)$, $g(t)$ Drift and diffusion coefficients, $f(t)=\tfrac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}$, $g(t)$ satisfies $g^{2}(t)=\tfrac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\tfrac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}$.
$\boldsymbol{f}(\boldsymbol{x}_{t})$ Transformation: $\tfrac{\boldsymbol{x}_{t}}{\alpha_{t}}$ (noise-prediction) or $\tfrac{\boldsymbol{x}_{t}}{\sigma_{t}}$ (data-prediction).
$\kappa(t),\psi(\tau)$ Time reparameterization $\kappa(t)$ and its inverse $\psi(\tau)$ with $\psi(\kappa(t))=t$.
$\boldsymbol{\iota}(\boldsymbol{x}_{t_{i-1}})$ Difference term: $\boldsymbol{f}(\boldsymbol{x}_{t_{i-1}})-\boldsymbol{f}(\boldsymbol{x}_{t_{i}})$.
$h_{t_{i}},\hat{h}_{t_{i}}$ Step sizes: $\kappa(t_{i-1})-\kappa(t_{i})$, auxiliary $\kappa(s_{i})-\kappa(t_{i})$.
$h_{\lambda_{i}}$ Step size in the log-SNR space: $\lambda(t_{i-1})-\lambda(t_{i})$.
$r_{i},r_{\text{logSNR}}(i)$ Ratio of consecutive step sizes in log-SNR space, used for gradient estimation.
$F_{\theta},B_{\theta},\bar{B}_{\theta}$ Finite difference terms for gradient approximation (Forward, Backward, and unified).
$\mathrm{Var}(\cdot|\cdot),\displaystyle H_{p}(\cdot|\cdot),\boldsymbol{\mu}_{t_{i}|t_{i+1}},\boldsymbol{\Sigma}_{t_{i}}$ Conditional variance, entropy, mean, and covariance matrix.
$\zeta_{i},\bar{\zeta}_{i},\eta_{i},\mu$ Optimization parameters (interpolation, complement, balance, shift).
$\mathcal{L}(\theta),\mathcal{L}_{1},\mathcal{L}_{2},w(t)$ General training objective (e.g., Eq. (3)), specific optimization objectives ($\mathcal{L}_{1},\mathcal{L}_{2}$), and training weight $w(t)$.
$G(\zeta_{i}),P(\boldsymbol{x}_{t_{i-1}})$ Auxiliary interpolation and optimization terms.
$D_{i},E(t_{i-1},t_{i}),\tilde{\Delta}^{g}_{t_{i}}$ Difference, gradient error, and weighted gradient terms.
$\nabla_{\boldsymbol{x}},\mathbb{E}[\cdot]$ Gradient, expectation.
$\text{Tr}(\cdot),\det(\cdot),\|\cdot\|$ Trace, determinant, and norm operators.
$\text{vec}(\cdot)$ Vectorization operator.
$\mathbb{R}^{d},\boldsymbol{I},\mathbf{0},C$ $d$-dimensional real space, identity matrix, zero vector, Gaussian entropy constant $C=\tfrac{1}{2}d(\log 2\pi+1)$.